# Mpemba meets Newton: Exploring the Mpemba and Kovacs effects in the time-delayed cooling law

### v1

@@ II.1 General solution in Laplace space
$T(t)=\begin{cases}T_{\text{b}}^{-},&t<-t_{\text{w}},\\ T_{\text{b}}^{-*}+(T_{\text{b}}^{-}-T_{\text{b}}^{-*})\mathcal{E}(t+t_{\text{w}}),&-t_{\text{w}}<t<0,\\ T_{\text{b}}^{+}+(T_{\text{b}}^{-*}-T_{\text{b}}^{+})\mathcal{E}(t)\\ +(T_{\text{b}}^{-}-T_{\text{b}}^{-*})\mathcal{E}(t+t_{\text{w}}),&t>0.\end{cases}$ (22)
@@ III Mpemba effect
To explore whether the Mpemba effect can be observed from the solutions of the time-delayed Newton’s cooling equation, Eq. (3), we consider two thermal reservoirs, a hot one (at temperature $T_{\text{b}}^{\text{h}}$) and a cold one (at temperature $T_{\text{b}}^{\text{c}}$). Two simple protocols (A and B) are proposed for the Mpemba effect, both direct and inverse, as illustrated in Fig. 6. Let us first describe the protocols for the direct effect, Fig. 6(a). According to protocol A, one sample (A) is first equilibrated at the hot bath temperature $T_{\text{b}}^{\text{h}}$ and then it is suddenly quenched to the cold bath temperature $T_{\text{b}}^{\text{c}}$ at time $t=-t_{\text{w}}$. The other sample (B) is first equilibrated at the cold bath temperature $T_{\text{b}}^{\text{c}}$, then quenched to the hot bath temperature $T_{\text{b}}^{\text{h}}$ at $t=-t_{\text{w}}$, and finally quenched to the cold bath at $t=0$. Thus, sample A experiences a single quench ($T_{\text{b}}^{\text{h}}\to T_{\text{b}}^{\text{c}}$), while sample B is subject to a double quench ($T_{\text{b}}^{\text{c}}\to T_{\text{b}}^{\text{h}}\to T_{\text{b}}^{\text{c}}$). As seen from Fig. 6(b), protocols A and B for the inverse Mpemba effect are the same, except for the exchange $T_{\text{b}}^{\text{h}}\leftrightarrow T_{\text{b}}^{\text{c}}$ (i.e., A: $T_{\text{b}}^{\text{c}}\to T_{\text{b}}^{\text{h}}$; B: $T_{\text{b}}^{\text{h}}\to T_{\text{b}}^{\text{c}}\to T_{\text{b}}^{\text{h}}$).
@@
$T_{\text{A}}(t)=\begin{cases}T_{\text{b}}^{\text{h}},&t<-t_{\text{w}},\\ T_{\text{b}}^{\text{c}}+(T_{\text{b}}^{\text{h}}-T_{\text{b}}^{\text{c}})\mathcal{E}(t+t_{\text{w}}),&t>-t_{\text{w}},\end{cases}$ (24a)
$T_{\text{B}}(t)=\begin{cases}T_{\text{b}}^{\text{c}},&t<-t_{\text{w}},\\ T_{\text{b}}^{\text{h}}-(T_{\text{b}}^{\text{h}}-T_{\text{b}}^{\text{c}})\mathcal{E}(t+t_{\text{w}}),&-t_{\text{w}}<t<0,\\ T_{\text{b}}^{\text{c}}+(T_{\text{b}}^{\text{h}}-T_{\text{b}}^{\text{c}})\left[\mathcal{E}(t)-\mathcal{E}(t+t_{\text{w}})\right],&t>0.\end{cases}$ (24b)
Figure 7: Plot of $T_{\text{A}}(t)$ and $T_{\text{B}}(t)$, including their past history for $t<0$. Here, $T_{\text{b}}^{\text{h}}=2$, $T_{\text{b}}^{\text{c}}=1$, and $\tau=0.36$, with (a) $t_{\text{w}}=0.2$, (b) $t_{\text{w}}=0.4$, and (c) $t_{\text{w}}=0.6$. In each panel, the dashed line represents $T_{\text{B}}(t)$ were not the sample be quenched to $T_{\text{b}}^{\text{c}}$ at $t=0$. Note that a Mpemba effect exists in the case of panel (b) ($t_{\text{w}}=0.4$).
As an example, Fig. 7 shows the time evolution of $T_{\text{A}}(t)$ and $T_{\text{B}}(t)$ for a hot bath at $T_{\text{b}}^{\text{h}}=2$, a cold bath at $T_{\text{b}}^{\text{c}}=1$, a time delay $\tau=0.36$, and three waiting times ($t_{\text{w}}=0.2,0.4,0.6$). As we can see, if the waiting time is too short (for instance, $t_{\text{w}}=0.2$), $T_{\text{B}}(0)$ is so far below $T_{\text{A}}(0)$ that no Mpemba crossing is possible. On the other hand, if the waiting time is too long (for instance, $t_{\text{w}}=0.6$), a crossing takes place during the preparation stage, that is, before sample B is quenched to $T_{\text{b}}^{\text{c}}$ at $t=0$, so that $T_{\text{B}}(0)>T_{\text{A}}(0)$ and again no Mpemba effect occurs. However, if the waiting time is within the right range (for instance, $t_{\text{w}}=0.4$), one has $T_{\text{A}}(0)>T_{\text{B}}(0)$ but sample A cools down sufficiently faster than sample B as to eventually overtake it at a certain crossover time $t_{\times}$ (direct Mpemba effect).
[…]
Let us now investigate the necessary and sufficient conditions for the existence of the Mpemba effect. From Eqs. (III), we have, for $t\geq 0$,
@@
For the protocols shown in Fig. 6(b) one would have $T_{\text{A}}(t)-T_{\text{B}}(t)=-(T_{\text{b}}^{\text{h}}-T_{\text{b}}^{\text{c}})\Delta(t)$. Thus, our first observation is that the existence or absence of the Mpemba effect (both direct and inverse) is independent of the bath temperatures $T_{\text{b}}^{\text{h}}$ and $T_{\text{b}}^{\text{c}}$, depending only on the delay time $\tau$ and the waiting time $t_{\text{w}}$ through the difference function $\Delta(t)$. Figure 8 shows the function $\Delta(t)$ for the same values of $\tau$ and $t_{\text{w}}$ as in Fig. 7.
@@
Next, once $t_{\text{w}}<t_{\text{w}}^{\max}$ and, therefore, $\Delta(t)>0$, the occurrence of a Mpemba effect requires that $\Delta(t_{\times})=0$ at a certain crossover time $t_{\times}$, so that it asymptotically relaxes to $\Delta(t)\to 0$ from below, as exemplified by the curve with $t_{\text{w}}=0.4$ in Fig. 8. After the crossover time $t_{\times}$, $-\Delta(t)$ presents a maximum at a time $t_{\text{M}}$ given by the solution to $\dot{\Delta}(t)=0$; according to Eqs. (16) and (26), $\dot{\Delta}(t)=-\Delta(t-\tau)$, implying that $t_{\text{M}}=t_{\times}+\tau$.
From Eq. (17), se see that the asymptotic long-time behavior of $\Delta(t)$ is
The amplitude $A_{\Delta}$ is plotted in Fig. 9 as a function of $t_{\text{w}}/t_{\text{w}}^{\text{max}}$ for $\tau=0$, $0.3$, and $0.36$. Except in the undelayed case ($\tau=0$), we can see that $A_{\Delta}$ becomes negative if $t_{\text{w}}$ is larger than a threshold value $t_{\text{w}}^{\min}$. Therefore, given a time delay $\tau$, the minimum waiting time $t_{\text{w}}^{\min}$ is determined by the condition $A_{\Delta}=0$, that is, $t_{\text{w}}^{\min}=\kappa^{-1}\ln 2$.
[…]
Figure 11: Phase space for the Mpemba effect. The lower and upper curves represent $t_{\text{w}}^{\min}$ and $t_{\text{w}}^{\max}$, respectively, as functions of the delay time $\tau$. If $t_{\text{w}}<t_{\text{w}}^{\min}$, then $\Delta(t)>0$ for $t\geq 0$. If, on the other hand, $t_{\text{w}}>t_{\text{w}}^{\max}$, then $\Delta(t)<0$ for $t\geq 0$. Therefore, the Mpemba effect occurs if and only if $t_{\text{w}}^{\min}<t_{\text{w}}<t_{\text{w}}^{\max}$ (shaded region). The dashed line represents the locus $\Delta(0)=|\Delta(t_{\text{M}})|$.
Figure 12: Plot of $T_{\text{A}}(t)$ and $T_{\text{B}}(t)$ for $t>0$. Panels (a, d), (b, e), and (c, f) correspond to $(\tau,t_{\text{w}})=(0.25,0.52)$, $(\tau,t_{\text{w}})=(0.30,0.49)$, and $(\tau,t_{\text{w}})=(0.36,0.47)$, respectively. Panels (a–c) illustrate the direct Mpemba effect (with $T_{\text{b}}^{\text{c}}=1$ and $T_{\text{b}}^{\text{h}}=2$), whereas panels (d–f) illustrate the inverse Mpemba effect (with $T_{\text{b}}^{\text{c}}=0.5$ and $T_{\text{b}}^{\text{h}}=1$).
If $t_{\text{w}}^{\min}<t_{\text{w}}<t_{\text{w}}^{\text{max}}$, the crossover time $t_{\times}$ is characterized by the condition $\Delta(t_{\times})=0$. It is plotted in Fig. 10 as a function of $t_{\text{w}}/t_{\text{w}}^{\min}$ for $\tau=0.25$, $0.3$, and $0.36$. As $t_{\text{w}}$ increases from $t_{\text{w}}^{\min}$ to $t_{\text{w}}^{\max}$, the crossover time $t_{\times}$ decreases monotonically, vanishing at $t_{\text{w}}=t_{\text{w}}^{\max}$. Near $t_{\text{w}}=t_{\text{w}}^{\min}$, $t_{\times}$ diverges logarithmically as $t_{\times}\sim-\ln(t_{\text{w}}/t_{\text{w}}^{\min}-1)$ (see inset of Fig. 10).
In summary, the time-delayed Newton’s cooling equation exhibits a Mpemba effect (either direct or inverse) under the protocols depicted in Fig. 6 if and only if the two only control parameters ($\tau$ and $t_{\text{w}}$) are such that the difference function $\Delta(t)$ fulfills two conditions: (i) $\Delta(0)>0$ (implying $t_{\text{w}}<t_{\text{w}}^{\max}$) and (ii) $\Delta(t)<0$ in the long-time regime (implying $t_{\text{w}}>t_{\text{w}}^{\min}$). The phase space for the occurrence of the Mpemba effect on the plane $t_{\text{w}}$ vs $\tau$ is shown in Fig. 11. Within the shaded Mpemba region, one can say that, for a given time delay $\tau$, the magnitude of the effect is maximal if the waiting time $t_{\text{w}}$ is such that the maximum positive value and the minimum negative value of $\Delta(t)$ are the same (except for the sign), i.e., $\Delta(0)=|\Delta(t_{\text{M}})|$. That line of maximal Mpemba effect is also included in Fig. 11.
Some illustrations of the Mpemba effect (both direct and inverse) are displayed in Fig. 12. The chosen values of $\tau$ and $t_{\text{w}}$ are close to the locus of maximal effect. We observe that, as expected, the Mpemba effect tends to become more pronounced as the time delay $\tau$ increases.
@@
Figure 14: Kovacs effect for $\tau=0.36$ and $t_{\text{w}}=0.1$, $0.2$, …, $1$. Panel (a) corresponds to the direct effect (with $T_{\text{b}}^{\text{h}}=2$ and $T_{\text{b}}^{\text{c}}=1$), whereas panel (b) corresponds to the inverse effect (with $T_{\text{b}}^{\text{c}}=0.5$ and $T_{\text{b}}^{\text{h}}=1$). In each panel, the dashed line represents $T(t)$ were not the sample be quenched to $T(t_{\text{w}})$ at $t=t_{\text{w}}$.
@@ IV Kovacs effect
In the Kovacs effect, the protocol is similar to that of sample B in the Mpemba effect (see Fig. 6), except for a couple of points, as summarized in Fig. 13. First, it is convenient to shift time an amount $t_{\text{w}}$, so that the first quench occurs at $t=0$ and the second one at the waiting time $t=t_{\text{w}}$. Second, the final bath temperature is made to coincide with the instantaneous system’s temperature at $t=t_{\text{w}}$, i.e., $T_{\text{b}}^{+}=T(t_{\text{w}})\equiv T_{\text{b}}^{\text{w}}$. Thus, according to Eq. (22), the temperature in the direct Kovacs effect is
$T(t)=\begin{cases}T_{\text{b}}^{\text{h}},&t<0,\\ T_{\text{b}}^{\text{c}}+(T_{\text{b}}^{\text{h}}-T_{\text{b}}^{\text{c}})\mathcal{E}(t),&0<t<t_{\text{w}},\\ T_{\text{b}}^{\text{w}}+(T_{\text{b}}^{\text{c}}-T_{\text{b}}^{\text{w}})\mathcal{E}(t-t_{\text{w}})\\ +(T_{\text{b}}^{\text{h}}-T_{\text{b}}^{\text{c}})\mathcal{E}(t),&t>t_{\text{w}},\end{cases}$ (28)
@@
Figure 14 illustrates the direct and inverse effects for $\tau=0.36$. As can be seen, after the quench at $t=t_{\text{w}}$, the temperature presents a local minimum (maximum) in the direct or cooling (inverse or heating) case. In the direct Kovacs effect, the temperature slope experiences a discontinuity from $\dot{T}(t_{\text{w}}^{-})=-\left[T(t_{\text{w}}-\tau)-T_{\text{b}}^{\text{c}}\right]$ to $\dot{T}(t_{\text{w}}^{+})=-\left[T(t_{\text{w}}-\tau)-T_{\text{b}}^{\text{w}}\right]$ but the sign is negative at both sides of $t=t_{\text{w}}$. Consequently, the Kovacs hump appears below $T=T_{\text{b}}^{\text{w}}$, thus qualifying as an anomalous Kovacs effect [94, 99]. An analogous conclusion holds in the inverse effect, where $\dot{T}(t_{\text{w}}^{-})=T_{\text{b}}^{\text{h}}-T(t_{\text{w}}-\tau)>\dot{T}(t_{\text{w}}^{+})=T_{\text{b}}^{\text{w}}-T(t_{\text{w}}-\tau)>0$ and the hump appears above $T=T_{\text{b}}^{\text{w}}$.
Let us characterize the Kovacs hump in more detail. From Eqs. (28) and (29), one finds that the temperature in the domain $t>t_{\text{w}}$ is
@@
is a semi-definite positive function, henceforth named the Kovacs hump function, which characterizes the relative strength of the Kovacs effect. It vanishes both at $t=0$ and in the limit $t\to\infty$. Note that $\dot{K}(t)=-\mathcal{E}(t-\tau)+\mathcal{E}(t-\tau+t_{\text{w}})/\mathcal{E}(t_{\text{w}})$, so that $\dot{K}(\tau)=0$, implying that $K(t)$ has a maximum value $K_{\max}=1-\tau-\mathcal{E}(\tau+t_{\text{w}})/\mathcal{E}(t_{\text{w}})$ at a time $t_{\text{K}}=\tau$. At a given time delay $\tau$, $K_{\max}$ increases monotonically with increasing waiting time $t_{\text{w}}$, reaching a finite value in the limit $t_{\text{w}}\to\infty$. From Eq. (17), we have $\lim_{t_{\text{w}}\to\infty}K_{\max}=1-\tau-e^{-\kappa\tau}=1-\tau-\kappa^{-1}$.
Figure 15(a) shows the Kovacs hump function $K(t)$ for a time delay $\tau=0.36$ and several values of the waiting time. The maximum value $K_{\max}$ is plotted in Fig. 15(b) as a function of $t_{\text{w}}$ for several values of $\tau$. As expected, the relative strength of the Kovas effect, as measured by $K_{\max}$, increases with increasing $\tau$ y $t_{\text{w}}$.

### v2

@@ II.2 General solution in Laplace space
$T(t)=\begin{cases}T_{\text{b}}^{-},&t\leq-t_{\text{w}},\\ T_{\text{b}}^{-*}+(T_{\text{b}}^{-}-T_{\text{b}}^{-*})\mathcal{E}(t+t_{\text{w}}),&-t_{\text{w}}\leq t\leq 0,\\ T_{\text{b}}^{+}+(T_{\text{b}}^{-*}-T_{\text{b}}^{+})\mathcal{E}(t)\\ +(T_{\text{b}}^{-}-T_{\text{b}}^{-*})\mathcal{E}(t+t_{\text{w}}),&t\geq 0.\end{cases}$ (27)
@@ III Mpemba effect
To explore whether the Mpemba effect can be observed from the solutions of the time-delayed Newton’s cooling equation, given by Eq. (3), we consider two thermal reservoirs, i.e., a hot one (at temperature $T_{\text{b}}^{\text{h}}$) and a cold one (at temperature $T_{\text{b}}^{\text{c}}$). Two simple protocols (A and B) are proposed for the Mpemba effect, both direct and inverse, as illustrated in Fig. 6. Let us first describe the protocols for the direct effect; see Fig. 6(a). According to protocol A, one sample (A) is first equilibrated at the hot bath temperature $T_{\text{b}}^{\text{h}}$ and then it is suddenly quenched to the cold bath temperature $T_{\text{b}}^{\text{c}}$ at time $t=-t_{\text{w}}$. The other sample (B) is first equilibrated at the cold bath temperature $T_{\text{b}}^{\text{c}}$, then quenched to the hot bath temperature $T_{\text{b}}^{\text{h}}$ at $t=-t_{\text{w}}$, and finally quenched to the cold bath at $t=0$. Thus, sample A experiences a single quench ($T_{\text{b}}^{\text{h}}\to T_{\text{b}}^{\text{c}}$), while sample B is subjected to a double quench ($T_{\text{b}}^{\text{c}}\to T_{\text{b}}^{\text{h}}\to T_{\text{b}}^{\text{c}}$). As seen from Fig. 6(b), protocols A and B for the inverse Mpemba effect are the same, except for the exchange $T_{\text{b}}^{\text{h}}\leftrightarrow T_{\text{b}}^{\text{c}}$ (i.e., A: $T_{\text{b}}^{\text{c}}\to T_{\text{b}}^{\text{h}}$; B: $T_{\text{b}}^{\text{h}}\to T_{\text{b}}^{\text{c}}\to T_{\text{b}}^{\text{h}}$).
@@
$T_{\text{A}}(t)=\begin{cases}T_{\text{b}}^{\text{h}},&t\leq-t_{\text{w}},\\ T_{\text{b}}^{\text{c}}+(T_{\text{b}}^{\text{h}}-T_{\text{b}}^{\text{c}})\mathcal{E}(t+t_{\text{w}}),&t\geq-t_{\text{w}},\end{cases}$ (29a)
$T_{\text{B}}(t)=\begin{cases}T_{\text{b}}^{\text{c}},&t\leq-t_{\text{w}},\\ T_{\text{b}}^{\text{h}}-(T_{\text{b}}^{\text{h}}-T_{\text{b}}^{\text{c}})\mathcal{E}(t+t_{\text{w}}),&-t_{\text{w}}\leq t\leq 0,\\ T_{\text{b}}^{\text{c}}+(T_{\text{b}}^{\text{h}}-T_{\text{b}}^{\text{c}})\left[\mathcal{E}(t)-\mathcal{E}(t+t_{\text{w}})\right],&t\geq 0.\end{cases}$ (29b)
Figure 7: Plot of $T_{\text{A}}(t)$ and $T_{\text{B}}(t)$, including their history for $t<0$. Here, $T_{\text{b}}^{\text{h}}=2$, $T_{\text{b}}^{\text{c}}=1$, and $\tau=0.36$, with (a) $t_{\text{w}}=0.2$, (b) $t_{\text{w}}=0.4$, and (c) $t_{\text{w}}=0.6$. In each panel, the dashed line depicts $T_{\text{B}}(t)$ if the sample were not quenched to $T_{\text{b}}^{\text{c}}$ at $t=0$. Note that a Mpemba effect exists in the case of (b) ($t_{\text{w}}=0.4$).
As an example, Fig. 7 shows the time evolution of $T_{\text{A}}(t)$ and $T_{\text{B}}(t)$ for a hot bath at $T_{\text{b}}^{\text{h}}=2$ (in arbitrary units), a cold bath at $T_{\text{b}}^{\text{c}}=1$, a delay time $\tau=0.36$, and three waiting times ($t_{\text{w}}=0.2,0.4,0.6$). As we can see, if the waiting time is too short (for instance, $t_{\text{w}}=0.2$), $T_{\text{B}}(0)$ is so far below $T_{\text{A}}(0)$ that no Mpemba crossing is possible. On the other hand, if the waiting time is too long (for instance, $t_{\text{w}}=0.6$), a crossing takes place during the preparation stage, that is, before sample B is quenched to $T_{\text{b}}^{\text{c}}$ at $t=0$, so that $T_{\text{B}}(0)>T_{\text{A}}(0)$ and again no Mpemba effect occurs. However, if the waiting time is within the right range (for instance, $t_{\text{w}}=0.4$), one has $T_{\text{A}}(0)>T_{\text{B}}(0)$ but sample A cools down sufficiently faster than sample B as to eventually overtake it at a certain crossover time $t_{\times}$ (direct Mpemba effect).
Figure 8: Plot of the difference function $\Delta(t)$ for $\tau=0.36$ and $t_{\text{w}}=0.2$ (solid curve), $0.4$ (dash-dotted curve), and $0.6$ (dashed curve). The circle in the curve corresponding to $t_{\text{w}}=0.4$ denotes the corresponding crossover time $t_{\times}$.
Let us now investigate the necessary and sufficient conditions for the existence of the Mpemba effect. From Eqs. (29), we have, for $t\geq 0$,
@@
For the protocols shown in Fig. 6(b), one would have $T_{\text{A}}(t)-T_{\text{B}}(t)=-(T_{\text{b}}^{\text{h}}-T_{\text{b}}^{\text{c}})\Delta(t)$. Thus, our first observation is that the existence or absence of the Mpemba effect (both direct and inverse) is independent of the bath temperatures $T_{\text{b}}^{\text{h}}$ and $T_{\text{b}}^{\text{c}}$, depending only on the delay time $\tau$ and the waiting time $t_{\text{w}}$ through the difference function $\Delta(t)$. Figure 8 shows the function $\Delta(t)$ for the same values of $\tau$ and $t_{\text{w}}$ as in Fig. 7.
@@
Figure 9: Plot of the amplitude $A_{\Delta}$ as a function of $t_{\text{w}}/t_{\text{w}}^{\max}$ for $\tau=0$ (solid curve), $0.30$ (dash-dotted curve), and $0.36$ (dashed curve). The circles in the curves corresponding to $\tau=0.30$ and $0.36$ denote the respective values of $t_{\text{w}}^{\min}/t_{\text{w}}^{\max}$.
Next, once $t_{\text{w}}<t_{\text{w}}^{\max}$ and, therefore, $\Delta(t)>0$, the occurrence of a Mpemba effect requires that $\Delta(t_{\times})=0$ at a certain crossover time $t_{\times}$, so that it asymptotically relaxes to $\Delta(t)\to 0$ from below, as exemplified by the curve with $t_{\text{w}}=0.4$ in Fig. 8. After the crossover time $t_{\times}$, $-\Delta(t)$ presents a maximum at a time $t_{\text{M}}$ given by the solution to $\dot{\Delta}(t)=0$. According to Eqs. (21) and (31),
$\dot{\Delta}(t)=-\begin{cases}1,&0\leq t\leq\max\{0,\tau-t_{\text{w}}\},\\ 2\mathcal{E}(t+t_{\text{w}}-\tau)-1,&\max\{0,\tau-t_{\text{w}}\}\leq t\leq\tau,\\ \Delta(t-\tau),&t\geq\tau,\end{cases}$ (32)
implying that $t_{\text{M}}=t_{\times}+\tau$
From Eq. (22), we see that the asymptotic long-time behavior of $\Delta(t)$ is
The amplitude $A_{\Delta}$ is plotted in Fig. 9 as a function of $t_{\text{w}}/t_{\text{w}}^{\text{max}}$ for $\tau=0$, $0.3$, and $0.36$. Except in the undelayed case ($\tau=0$), we can see that $A_{\Delta}$ becomes negative if $t_{\text{w}}$ is larger than a threshold value $t_{\text{w}}^{\min}$. Therefore, given a delay time $\tau$, the minimum waiting time $t_{\text{w}}^{\min}$ is determined by the condition $A_{\Delta}=0$, that is, $t_{\text{w}}^{\min}=\kappa^{-1}\ln 2$.
Figure 10: Plot of the crossover time $t_{\times}$ as a function of $t_{\text{w}}/t_{\text{w}}^{\min}$ for $\tau=0.25$ (solid curve), $0.30$ (dash-dotted curve), and $0.36$ (dashed curve). The circles denote the respective values of $t_{\text{w}}^{\max}/t_{\text{w}}^{\min}$. Inset: $t_{\times}$ vs $\log_{10}(t_{\text{w}}/t_{\text{w}}^{\min}-1)$.
Figure 11: Phase space for the Mpemba effect. The lower and upper curves represent $t_{\text{w}}^{\min}$ and $t_{\text{w}}^{\max}$, respectively, as functions of the delay time $\tau$. If $t_{\text{w}}<t_{\text{w}}^{\min}$, then $\Delta(t)>0$ for $t\geq 0$. If, on the other hand, $t_{\text{w}}>t_{\text{w}}^{\max}$, then $\Delta(t)<0$ for $t\geq 0$. Therefore, the Mpemba effect occurs if and only if $t_{\text{w}}^{\min}<t_{\text{w}}<t_{\text{w}}^{\max}$ (shaded region). The dashed line represents the locus defined by the condition $\Delta(0)=|\Delta(t_{\text{M}})|$.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 12: Plot of $T_{\text{A}}(t)$ and $T_{\text{B}}(t)$ for $t>0$. (a),(d) $(\tau,t_{\text{w}})=(0.25,0.52)$, (b),(e) $(\tau,t_{\text{w}})=(0.30,0.49)$, and (c),(f) $(\tau,t_{\text{w}})=(0.36,0.47)$. (a)–(c) The direct Mpemba effect (with $T_{\text{b}}^{\text{c}}=1$ and $T_{\text{b}}^{\text{h}}=2$); (d)–(f) the inverse Mpemba effect (with $T_{\text{b}}^{\text{c}}=0.5$ and $T_{\text{b}}^{\text{h}}=1$).
If $t_{\text{w}}^{\min}<t_{\text{w}}<t_{\text{w}}^{\text{max}}$, the crossover time $t_{\times}$ is characterized by the condition $\Delta(t_{\times})=0$. It is plotted in Fig. 10 as a function of $t_{\text{w}}/t_{\text{w}}^{\min}$ for $\tau=0.25$, $0.3$, and $0.36$. As $t_{\text{w}}$ increases from $t_{\text{w}}^{\min}$ to $t_{\text{w}}^{\max}$, the crossover time $t_{\times}$ decreases monotonically, vanishing at $t_{\text{w}}=t_{\text{w}}^{\max}$. Near $t_{\text{w}}=t_{\text{w}}^{\min}$, $t_{\times}$ diverges logarithmically as $t_{\times}\sim-\log_{10}(t_{\text{w}}/t_{\text{w}}^{\min}-1)$ (see inset of Fig. 10).
@@
Some illustrations of the Mpemba effect (both direct and inverse) are displayed in Fig. 12. The chosen values of $\tau$ and $t_{\text{w}}$ are close to the locus of maximal effect. We observe that, as expected, the Mpemba effect tends to become more pronounced as the delay time $\tau$ increases.
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Figure 14: Kovacs effect for $\tau=0.36$ and $t_{\text{w}}=0.1$, $0.2$, …, $1$. (a) The direct effect (with $T_{\text{b}}^{\text{h}}=2$ and $T_{\text{b}}^{\text{c}}=1$); (b) the inverse effect (with $T_{\text{b}}^{\text{c}}=0.5$ and $T_{\text{b}}^{\text{h}}=1$). In each panel, the dashed line represents $T(t)$ if the sample were not quenched to $T(t_{\text{w}})$ at $t=t_{\text{w}}$.
Refer to captionRefer to caption
@@ IV Kovacs effect
In the Kovacs effect, the protocol is similar to that of sample B in the Mpemba effect (see Fig. 6), except for a couple of points, as summarized in Fig. 13. First, it is convenient to shift time by an amount $t_{\text{w}}$, so that the first quench occurs at $t=0$ and the second one at a waiting time $t=t_{\text{w}}$. Second, the final bath temperature is made to coincide with the instantaneous system’s temperature at $t=t_{\text{w}}$, i.e., $T_{\text{b}}^{+}=T(t_{\text{w}})\equiv T_{\text{b}}^{\text{w}}$. Thus, according to Eq. (27), the temperature in the direct Kovacs effect is
$T(t)=\begin{cases}T_{\text{b}}^{\text{h}},&t\leq 0,\\ T_{\text{b}}^{\text{c}}+(T_{\text{b}}^{\text{h}}-T_{\text{b}}^{\text{c}})\mathcal{E}(t),&0\leq t\leq t_{\text{w}},\\ T_{\text{b}}^{\text{w}}+(T_{\text{b}}^{\text{c}}-T_{\text{b}}^{\text{w}})\mathcal{E}(t-t_{\text{w}})\\ +(T_{\text{b}}^{\text{h}}-T_{\text{b}}^{\text{c}})\mathcal{E}(t),&t\geq t_{\text{w}},\end{cases}$ (34)
@@
Figure 14 illustrates the direct and inverse effects for $\tau=0.36$. As can be seen, after the quench at $t=t_{\text{w}}$, the temperature presents a local minimum (maximum) in the direct or cooling (inverse or heating) case. In the direct Kovacs effect, the temperature slope experiences a discontinuity from $\dot{T}(t_{\text{w}}^{-})=-\left[T(t_{\text{w}}-\tau)-T_{\text{b}}^{\text{c}}\right]$ to $\dot{T}(t_{\text{w}}^{+})=-\left[T(t_{\text{w}}-\tau)-T_{\text{b}}^{\text{w}}\right]$, but the sign is negative at both sides of $t=t_{\text{w}}$. Consequently, the Kovacs hump appears below $T=T_{\text{b}}^{\text{w}}$, thus qualifying as an anomalous Kovacs effect [108, 114]. An analogous conclusion holds in the inverse effect, where $\dot{T}(t_{\text{w}}^{-})=T_{\text{b}}^{\text{h}}-T(t_{\text{w}}-\tau)>\dot{T}(t_{\text{w}}^{+})=T_{\text{b}}^{\text{w}}-T(t_{\text{w}}-\tau)>0$ and the hump appears above $T=T_{\text{b}}^{\text{w}}$.
Let us characterize the Kovacs hump in more detail. From Eqs. (34) and (35), one finds that the temperature in the domain $t\geq t_{\text{w}}$ is
@@
is a semi-definite positive function, henceforth named the Kovacs hump function, which characterizes the relative strength of the Kovacs effect. It vanishes both at $t=0$ and in the limit $t\to\infty$. Note that
$\dot{K}(t)=-\begin{cases}1-\frac{1}{\mathcal{E}(t_{\text{w}})},&0\leq t\leq\max\{0,\tau-t_{\text{w}}\},\\ 1-\frac{\mathcal{E}(t+t_{\text{w}}-\tau)}{\mathcal{E}(t_{\text{w}})},&\max\{0,\tau-t_{\text{w}}\}\leq t\leq\tau,\\ K(t-\tau),&t\geq\tau.\end{cases}$ (38)
Therefore, $\dot{K}(\tau)=0$, implying that $K(t)$ has a maximum value $K_{\max}=1-\tau-\mathcal{E}(\tau+t_{\text{w}})/\mathcal{E}(t_{\text{w}})$ at a time $t_{\text{K}}=\tau$. At a given delay time $\tau$, $K_{\max}$ increases monotonically with increasing waiting time $t_{\text{w}}$, reaching a finite value in the limit $t_{\text{w}}\to\infty$. From Eq. (22), we have $\lim_{t_{\text{w}}\to\infty}K_{\max}=1-\tau-e^{-\kappa\tau}=1-\tau-\kappa^{-1}$.
Figure 15(a) shows the Kovacs hump function $K(t)$ for a delay time $\tau=0.36$ and several values of the waiting time. The maximum value $K_{\max}$ is plotted in Fig. 15(b) as a function of $t_{\text{w}}$ for several values of $\tau$. As expected, the relative strength of the Kovacs effect, as measured by $K_{\max}$, increases with increasing $\tau$ and $t_{\text{w}}$.